OpenStudy (anonymous):

A glass cylinder is completely filled with 32m6n2 cubic inches of salt solution. The jug has a base area of 8m3n2 square inches. The height of the jug, in inches, is represented by the expression The quantity of 32 m to the sixth n squared all over 8 m cubed n squared. Which of the following simplified expressions represents the height of the jug in inches? 4m9n4 4m2n 4m2 4m3

jimthompson5910 (jim_thompson5910):

Volume = (Area of base)*(height)

jimthompson5910 (jim_thompson5910):

wait a second...are there 2 figures? A cylinder and a jug?

OpenStudy (anonymous):

idk I'm guessing it's just one.

jimthompson5910 (jim_thompson5910):

oh i see now, nvm

jimthompson5910 (jim_thompson5910):

they're saying " The height of the jug, in inches, is represented by the expression The quantity of 32 m to the sixth n squared all over 8 m cubed n squared." so.. \[\Large h = \frac{32m^6n^2}{8m^3n^2}\]

jimthompson5910 (jim_thompson5910):

reduce that fraction to get your answer

OpenStudy (anonymous):

can you reduce it for me please?

jimthompson5910 (jim_thompson5910):

let's break it up

jimthompson5910 (jim_thompson5910):

32/8 = ???

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

m^6 divided by m^3 = ??? (subtract the exponents when dividing expressions like this with the same base)

OpenStudy (anonymous):

m3

jimthompson5910 (jim_thompson5910):

m^3, yep

jimthompson5910 (jim_thompson5910):

n^2 divided by n^2 = ???

OpenStudy (anonymous):

it would be 0 so the n wouldn't be there anymore

jimthompson5910 (jim_thompson5910):

we'd get n^0

jimthompson5910 (jim_thompson5910):

so, \[\Large h = \frac{32m^6n^2}{8m^3n^2}\] \[\Large h = 4m^3n^0\] \[\Large h = 4m^3*1\] \[\Large h = 4m^3\]

jimthompson5910 (jim_thompson5910):

keep in mind that anything to the 0th power is 1 (except 0 itself) so, \[\Large n^0 = 1\]

OpenStudy (anonymous):

so the final answer is \[4m^{3}\]

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome